OpenStudy (anonymous):

r=the squarer root of 2A/0 (the zero has a line in the middle idk what that is)]

OpenStudy (saifoo.khan):

Use the draw tool to make it.

OpenStudy (anonymous):

r=\[\sqrt{?}\]2a/\[\theta\]

OpenStudy (anonymous):

10\[\le\]4x-2\[\le\]22